Calculate the pressure in a 212 Liter tank containing 23.3 kg of argon gas at 25°C? ( atomic weight Ar - 40 )


1
Expert's answer
2022-06-06T17:01:05-0400

V = 212 l

T = 25°C =298 K

M(Ar) = 40 g/mol

R = 8,31

n(Ar) = m / M = 23300 / 40= 582,5 mol


P = nRT / V = 582,5 *8,31*298/212= 6804,2


Answer: 6804,2 kPa